OpenStudy (anonymous):

The population of a city increases by 4,000 people each year. In 2025, the population is projected to be 450,000 people. What is an equation that gives the city’s population p (in thousands of people) x years after 2010?

OpenStudy (anonymous):

450,000 = y + 4000* 15 where y is population in 2010 y = 450,000 - 60,000 = 390,000 required equation is p = 390,000 + 4000x